Can someone help me solve this?
Solve the system by the elimination method.
x/3+3y/4=-3
x/3-y/4=5

Answer by stanbon(75887)

Solve the system by the elimination method.
x/3+3y/4=-3
x/3-y/4=5
-----
Subtract and solve for "y"::
y = -8
--------
Solve for "x":::
x/3 - (1/4)y = 5
x/3 + (1/4)8 = 5
x/3 + 2 = 5
x/3 = 3
x = 9
================
Cheers,
Stan H.
================